# The AI Community Building the Future? A Quantitative Analysis of Development Activity on Hugging Face Hub

### v1

@@ 5.2.1 Recommendations for Policymakers
5.2.2 Recommendations for Platform Providers
We make two recommendations to HF as a platform for open model development. First, HF could work with researchers to identify features and API improvements that would aid research efforts concerning open model development on its platform, building on efforts by members of the HF community, such as Weyaxi/huggingface-leaderboard. This collaboration could include collecting and publishing data on open model development patterns and collaboration, which would help fill the current “data gap” in this area. HF may take inspiration from GitHub’s Innovation Graph [129] or its annual Octoverse reports [130], which provide access to data and insights on development activity on its platform. Second, a concerning proportion of models (64.67%) and datasets (72.13%) lack licences, which may be due to uncertainty about how or whether they should be licenced [131, 65]. For comparison, the number of unlicenced repositories on GitHub is lower at 46% or 53% if including “other licences” [132]. In the interest of promoting responsible development, HF should consider developing educational resources on licences, such as guides or tutorial videos, or developing features, such as a licence drop-down menu, which can inform developers of the options available as well as their merits and drawbacks. Such a feature could be considered amongst other recommendations to moderate models on the HF Hub, such as hiring AI safety researchers and proactively red-teaming unsafe models [52, 68].
@@ Acknowledgements
The authors would like to thank Loubna Ben Allal, Daniel van Strien, Peter Cihon, Mer Joyce, Stefano Maffulli, Matt White, Seb Elmes, David Gray Widder, and Alek Tarkowski as well as the peer-reviewers at the Journal of Computational Social Science for their generous feedback on prior versions of this manuscript.
@@
Appendix A Definition of Network Properties for Social Network Analysis
@@ Appendix C Mann-Whitney U Tests for Activity in Model Repositories
Restrictive No licence No licence

### v2

@@ 5.2 Implications for Practice
5.2.1 Recommendations for Open Source Practitioners
Beyond our academic research suggestions, we encourage open source researchers and practitioners to develop standardised metrics for studying open model development. Groups like the Linux Foundation’s Community Health Analytics in Open Source Software (CHAOSS) working group [133], which has created metrics to assess the health and sustainability of OSS developer communities, are well-positioned to lead this effort. The lack of empirical data on open model development hinders evidence-based decision-making in this rapidly evolving field, and by working together to establish appropriate metrics, open source practitioners can help to address the data gap in “open source AI”.
5.2.2 Recommendations for Platform Providers
We make two recommendations to HF as a platform for open model development. First, HF could work with researchers to identify features and API improvements that would aid research efforts concerning open model development on its platform, building on efforts by members of the HF community, such as Weyaxi/huggingface-leaderboard. This collaboration could include collecting and publishing data on open model development patterns and collaboration, which would help fill the current “data gap” in this area. HF may take inspiration from GitHub’s Innovation Graph [134] or its annual Octoverse reports [135], which provide access to data and insights on development activity on its platform. Second, a concerning proportion of models (64.67%) and datasets (72.13%) lack licenses, which may be due to uncertainty about how or whether they should be licensed [136, 67]. For comparison, the number of unlicensed repositories on GitHub is lower at 46% or 53% if including “other licenses” [137]. In the interest of promoting responsible development, HF should consider developing educational resources on licenses, such as guides or tutorial videos, or developing features, such as a license drop-down menu, which can inform developers of the options available as well as their merits and drawbacks. Such a feature could be considered amongst other recommendations to moderate models on the HF Hub, such as hiring AI safety researchers and proactively red-teaming unsafe models [53, 70].
@@ Acknowledgements
The authors would like to thank Loubna Ben Allal, Daniel van Strien, Peter Cihon, Mer Joyce, Stefano Maffulli, Matt White, Seb Elmes, David Gray Widder, Alek Tarkowski, Johan Linåker, Sean P. Goggins, and the reviewers at the Journal of Computational Social Science for their generous feedback on prior versions of this manuscript.
@@
Appendix A Definitions of Network Properties in Social Network Analysis
@@ Appendix C Mann-Whitney U Tests for Activity in Model Repositories
Restrictive No license No license